# references.bib

### v2

@@ Acknowledgement
9:     Update jump clocks, $\tilde{J}_{k}$, to $t+\Delta t$ ($\triangleright$ As per Equation (S11))
@@
12:        Compute time of jump event, $\Delta\tau$ ($\triangleright$ As per Equation (S12))
@@
The Operator Splitting Jump-Switch-Flow sampler is described in Algorithm S2. Unlike the Exact JSF sampler, when a jump event occurs, we continue to integrate the system with the same flow event, $\Delta\vec{V}_{F}(t)$, instead of resampling it as before. This, in turn, ensures that the ODE is not continually being resampled. While this adds complexity at the implementation stage, significant computational time saving is achieved.
@@
Algorithm S2 An algorithmic description of the Operator Splitting Jump-Switch-Flow sampler .
0:  $\vec{V}(0)$, $\Delta t>0$ and $T_{\max}>0$
@@
S3 Simulation study: SIRS with demography
@@
We first present how the SIRS model with demography can be expressed with the Jump-Switch-Flow method. For the Jump-Switch-Flow method to be a useful method, it must reproduce behaviour that is representative of the exact process (i.e. the CTMC). That is, given some summary statistic, distributions produced via sampling the JSF process should be comparable to those obtained via sampling the CTMC. Moreover, we require that it does so with acceptable computational efficiency. We show, through simulation experiments, how the Jump-Switch-Flow method compares to the Doob-Gillespie method (gillespie1976general) for accuracy, and how our approach can exhibit sensitivity with respect to model inputs. We present how our Jump-Switch-Flow method compares to both the Doob-Gillespie and Tau-Leaping methods.
@@
Figure 2(c) shows the SIRS model with demography as a Jump-Switch-Flow process. In this figure, we assume that the $S$ and $R$ compartments are continuous (flowing), while the $I$ compartment is discrete (jumping). To decide which of the arrows/reactions are jumping, i.e. in $\mathcal{S}$, and which are flowing, i.e. in $\mathcal{S}^{\prime}$, we consider the reactants and products of the arrows: if an arrow has any discrete reactant or product, then that arrow is also jumping, otherwise it is flowing. Table S1 shows the reactants and products of the SIRS model with demography.
Figure S2: 2(a) The labelled compartmental SIRS model with demography. 2(b) The SIRS model with demography where the arrows describe the interactions between the reactants and products. Here, blue arrows do not result in any flow through the system, and can be ignored. 2(c) The model as a Jump-Switch-Flow system with continuous (flowing) S and R (black), and discrete (jumping) I (purple). Since I is jumping, the reactions involving I are modelled as discrete, jumping reactions.
@@
Table S1: The stoichiometric matrices for the SIRS model with demography. The (*) indicates reactions that are jumping when $I$ is discrete and $S$ and $R$ are continuous. The rows with blue coloured entries do not alter the exchange through the system, and so need not be represented.
For the remainder of this section, we suppose parameter values for the SIRS model with demography that would be typical for a newly introduced, yearly seasonal communicable disease, as presented in Table S2. With these parameter values, on average, we expect an infected individual to infect 2 other people per week (during the initial outbreak), and be infectious for 1 week, with immunity to the disease lasting for 1 year. For the population turnover, we assume individuals live for, on average, 85 years. Initially there are two infectious individuals (i.e. $I(0)=2$), no recovered individuals (i.e. $R(0)=0$), and the remainder of the population is susceptible (i.e. $S(0)=N(0)-I(0)$).
@@
Value $2/7$ $1/7$ $1/365$ $1/(85\times 365)$ $1/(85\times 365)$
Table S2: Parameter values used for SIRS model with demography.
@@
    Extinction: all of the individuals recover before the disease manages to spread significantly throughout the population (see Figure 3(a) and 3(d));
@@
    Fade-out: the disease spreads throughout the population, however it fades-out in the trough proceeding the first wave (see Figure 3(b) and 3(e));
@@
    Endemic: the disease persists indefinitely within the population (see Figure 3(c) and 3(f)).
@@
Figure S3: Possible simulated trajectories of SIRS model with demography for an initial population size $N(0)=10^{5}$, from Doob-Gillespie (yellow) and Jump-Switch-Flow (purple, $\Omega=10^{3}$) methods: 3(a) two examples of extinction trajectories of the infectious compartment, with associated $S-I$ phase plane in 3(d); 3(b) two examples of fade-out trajectories of the infectious compartment, with associated $S-I$ phase plane in 3(e); 3(c) two examples of endemic trajectories of the infectious compartment, with associated $S-I$ phase plane in 3(f).
S3.1 Simulation experiments: Comparing Jump-Switch-Flow and Doob-Gillespie
We compare our Jump-Switch-Flow method to the gold standard approach for simulating exact solutions to CTMCs, the Doob-Gillespie method. To do so, we first specified an initial population size of $N(0)=10^{5}$, and a switching threshold of each compartment of $\Omega=10^{3}$. We then generated 5,000 simulations using the parameters in Table S2.
To capture the dynamics of the model, we track the infectious compartment as a key compartment of interest. In doing so, we then obtain distributions for the peak number of infectious individuals and also the time for the infection to peak, for the fade-out scenario. We also track the cumulative number of infected individuals after a certain amount of time for the endemic scenario. These results are presented in Figure S5.
@@
Figure S4: We compare the infectious compartment from realisations for Jump-Switch-Flow 4(b) and Doob-Gillespie 4(g) solutions, for a population size of $N(0)=10^{5}$. We use a switching threshold of $\Omega=10^{3}$. We also show the distributions of the time to infectious peak (4(a) and 4(f)), and infectious peak values (4(c) and 4(h)). We compare the cumulative infections over the simulations for Jump-Switch-Flow 4(d) and Doob-Gillespie 4(i) solutions and the final distributions (4(e) and 4(j)). Point averages are also presented as a solid line.
Comparing Figures 4(a) and 4(f), we can immediately observe that the distributions for time to peak for Jump-Switch-Flow method appears to be representative of that for the Doob-Gillespie method. This is because the delay in the epidemic take off at low population is fully captured with this sufficiently large choice of switching threshold, $\Omega=10^{3}$, enabling the full stochastic effects to be accurately represented.
However, if we compare the distributions for the peak number of infectious individuals (4(c) and 4(h)), we can see that they differ significantly, for this chosen switching threshold. This difference can be understood if we consider how the process is behaving. Once the infectious compartment is sufficiently large for the Jump-Switch-Flow method, the method switches from a stochastic representation to a deterministic representation. Therefore, once the infectious compartment becomes deterministic, given that the susceptible population is also deterministic, its peak value is determined by the number of both infectious and susceptible individuals at this time. Since the number of infectious people at this time will always be the same, $I=\Omega=10^{3}$ people, any variation in peak value is due to the number of susceptible individuals, $S$, at the time of switching.
If the summary statistic we instead compare is the distribution of the cumulative number of infected individuals after four years of epidemic circulation, we observe that, even for a relatively low switching threshold, the Jump-Switch-Flow method performs well comparatively to the Doob-Gillespie method, see Figures 4(e) and 4(j). We also note that for these parameter choices and switching threshold, the infectious compartment switches between jumping and flowing states multiple times, yet the Jump-Switch-Flow method produces a cumulative distribution comparable to that of Doob-Gillespie. This indicates that the Jump-Switch-Flow method is robust for long term simulations.
The final summary statistics we are interested in are the probabilities of extinction, fade-out and endemic scenarios. Table S3 shows the probability of each scenario computed from 5,000 simulated samples. Here, we see that both methods produce comparatively similar results, which indicates that the Jump-Switch-Flow method is capable of capturing the inherent stochasticity in the exact simulation of the system via the Doob-Gillespie method.
@@
Jump-Switch-Flow $0.2476\pm 0.0120$ $0.4774\pm 0.0138$ $0.2750\pm 0.0124$
Doob-Gillespie $0.2520\pm 0.0120$ $0.4668\pm 0.0138$ $0.2812\pm 0.0125$
Table S3: Probabilities (with associated 95% confidence interval) of scenario outcomes for Jump-Switch-Flow and Doob-Gillespie, from 5,000 simulations.
We also compare the Jump-Switch-Flow and Doob-Gillespie methods where the switching threshold has been set to the population size, resulting in no compartments switching into the flowing state. These results are further discussed in S3.2, where we show the two methods are indistinguishable.
S3.2 Simulation experiments: Comparing Jump-Switch-Flow (with no switching) and Doob-Gillespie
We compare the Jump-Switch-Flow method to the Doob-Gillespie method. To implement the latter, we set the switching threshold to be the total population size, to ensure no compartments switch into a flowing regime. We can therefore investigate how our Jumping process compares to exact solutions of the Continuous Time Markov Chains.
@@
Figure S5: We compare the infectious compartment from realisations for Jump-Switch-Flow 5(b) and Doob-Gillespie 5(g) solutions, for a population size of $N_{0}=10^{5}$. We use a switching threshold of $\Omega=10^{3}$. We also show the distributions of the time to infectious peak (5(a) and 5(f)), and infectious peak values (5(c) and 5(h)). We compare the cumulative infections over the simulations for Jump-Switch-Flow 5(d) and Doob-Gillespie 5(i) solutions and the final distributions (5(e) and 5(j)). Point averages are also presented as a solid line.
@@
Table S4 shows that the probabilities of scenario outcomes are also comparable for both Jump-Switch-Flow (with no switching) and Doob-Gillespie.
@@
Jump-Switch-Flow $0.2512\pm 0.0120$ $0.4700\pm 0.0139$ $0.2788\pm 0.0123$
Doob-Gillespie $0.2520\pm 0.0120$ $0.4668\pm 0.0139$ $0.2812\pm 0.0125$
Table S4: Probabilities (with associated 95% confidence interval) of scenario outcomes for Jump-Switch-Flow (with no switching) and Doob-Gillespie, from 5,000 simulations.
S3.3 Simulation experiments: Sensitivity of Jump-Switch-Flow to switching threshold
In this section, we demonstrate how the accuracy of the Jump-Switch-Flow method can exhibit sensitivity with respect to the switching threshold (Figure S6). In Figure 6(a), we can see that the distribution of peak number of infective individuals for the fade-out scenario is highly sensitive to the switching threshold. For a small switching threshold ($\Omega=10^{1}$), we see that the distribution exhibits low variance, is representative of the deterministic expectation, and is not capturing the distribution obtained with the Doob-Gillespie method. As the switching threshold is increased, we see that the distribution of the peak number of infective individuals increases in variance. However, the Jump-Switch-Flow method is not representative of the Doob-Gillespie method until a large switching threshold ($\Omega=N(0)=10^{5}$), where all compartments are in the jumping state.
Figure S6: Summary statistics of interest to observe the sensitivity of the Jump-Switch-Flow method with respect to the switching threshold. 6(a) Sensitivity to the distribution of peak number of infective individuals for the fade-out scenario. 6(c) Sensitivity to the distribution of peak infection time for the fade-out scenario. 6(b) Sensitivity to the distribution of the cumulative number of infected individuals after four years for the endemic scenario. 6(d) Sensitivity to the probability of each scenario occurring. In each of the plots 6(a)-6(c) grey to purple shaded distributions represent results obtained with the Jump-Switch-Flow method whilst the yellow distributions are the gold standard Doob-Gillespie algorithm (for which the placeholder ‘G’ is used to signify) which exhibit exact stochasticity but at the cost of high computational requirements.
@@
S3.4 Simulation experiments: Computational efficiency of Jump-Switch-Flow
@@
We now present the computational efficiency of the Jump-Switch-Flow method, and compare it to the Doob-Gillespie method and Tau-Leaping method (which is regarded as a computationally efficient, approximate, method).

### v3

@@ Acknowledgement
9:  Update jump clocks, $\tilde{J}_{k}$, to $t+\Delta t$ ($\triangleright$ As per Equation (16))
@@
12:   Compute time of jump event, $\Delta\tau$ ($\triangleright$ As per Equation (18))
@@
The Operator Splitting Jump-Switch-Flow sampler is described in Algorithm SI2. Unlike the Exact JSF sampler, when a jump event occurs, we continue to integrate the system with the same flow event, $\Delta\vec{V}_{F}(t)$, instead of resampling it as before. This, in turn, ensures that the ODE is not continually being resampled. While this adds complexity at the implementation stage, significant computational time saving is achieved.
@@
Algorithm SI2 An algorithmic description of the Operator Splitting Jump-Switch-Flow sampler .
$\vec{V}(0)$, $\Delta t>0$ and $T_{\max}>0$
@@
SI2 Simulation Study: Birth-death processes
Fundamentally, one can conceptualise any compartmental model as a collection of interacting birth-death processes. For this reason, we will consider two very simple such processes: 1) a birth dominated birth-death process, and 2) a death dominated birth-death process. In this section, we will perform a simulation study to highlight the advantages of our Jump-Switch-Flow method, and compare the results to the gold standard, Doob-Gillespie Exact method, and the highly computationally efficient Tau-hybrid method, which is provided by GillesPy2, written in both C++ and Python.
The exact model we will utilise is a single species, $X$, experiencing birth with rate $\alpha$, which results in $X\rightarrow 2X$, and death with rate $\beta$, which results in $X\rightarrow\varnothing$.
A birth dominated birth-death process
For the birth dominated process, we require $\alpha>\beta$. Therefore, we specify $\alpha=1$, and $\beta=0.5$. We also choose the initial species population $X(0)=1$, and a simulation time of $t=20$. The quantities of interest to test the accuracy and efficiency of the methods are (i) the CPU time for a particular simulation instance, (ii) the simulation time required to reach a particular value, $X=200$, and (iii) the accuracy of the method to experience initial population extinction. Figure SI1 shows results for each of the three quantities of interest for the Tau-hybrid method (red), Doo-Gillespie (yellow) and three different Jump-Switch-Flow examples (purple) with varying switching thresholds of $\Omega=10^{1}$, $\Omega=10^{2}$, and $\Omega=10^{3}$, respectively. Here, 1000 unique trajectories are considered for each method.
Figure SI1: Summary statistics for the birth dominated birth-death process for the Tau-hybrid (red) JSF for various switching thresholds, $\Omega$, (purple) and the Doob-Gillespie (yellow) methods. Left: CPU time required. We observe that as $\Omega$ increases, the computational time increases, but remains the computationally least expensive. Middle: Distributions in time to reach $X=200$. We observe that JSF results in comparable distributions to the exact process, with diminishing returns for $\Omega>10^{2}$. Right: Extinction probability. We observe that all point estimates are comparable to that of the exact method.
Considering the CPU time, we can observe as we increase the switching threshold, JSF requires more CPU time, since more of the simulation is in a discrete, Jumping regime. We also observe that the CPU time is, at worst, around two orders of magnitude less, compared to that of Doob-Gillespie. In comparison, we observe that the C++ implementation of Tau-hybrid method requires at best three times the computational time to simulate the equivalent process as JSF, while the Python implementation is at least one order of magnitude slower, indicating that JSF is much more efficient. To assess the accuracy of each method, we now compare the (simulation) time for the species to reach 200. Here, both Tau-hybrid and JSF with switching thresholds of $\Omega=10^{2}$, and $\Omega=10^{3}$ result in distributions and median value comparable to that of Doob-Gillespie. The only exception is that of JSF with a switching threshold of $\Omega=10^{1}$. Lastly, we observe that all methods result in an initial extinction probability comparable to that of Doob-Gillespie.
These results suggest that JSF is capable of recovering the key quantities of the exact stochastic method, while being computationally cheaper than current hybrid methods and the exact method, for the birth dominated birth-death process.
A death dominated birth-death process
For the birth dominated process, we require $\alpha<\beta$. Therefore, we specify $\alpha=0.5$, and $\beta=1$. We choose the initial species population $X(0)=100$, and a simulation time of $t=30$. As before, the quantities of interest to test the accuracy and efficiency of the methods are (i) the CPU time for a particular simulation instance, (ii) the simulation time required to reach a particular value, $X=50$, and (iii) the simulation time require for the species to go extinct. Figure SI2 shows results for each of the three quantities of interest for the Tau-hybrid method (red), Doo-Gillespie (yellow) and three different Jump-Switch-Flow examples (purple) with varying switching thresholds of $\Omega=10^{1}$, $\Omega=10^{2}$, and $\Omega=10^{3}$, respectively. Again, 1000 unique trajectories are considered for each method.
Figure SI2: Summary statistics for the death dominated birth-death process for the Tau-hybrid (red) JSF for various switching thresholds, $\Omega$, (purple) and the Doob-Gillespie (yellow) methods. Left: CPU time required. We again observe that as $\Omega$ increases, the computational time increases, but still remains the computationally least expensive. Middle: Distributions in time to reach $X=50$. We observe that a switching threshold of $\Omega=10^{3}$ is required for JSF to result in comparable distributions to the exact process, Right: Distribution in the time to extinction. We observe that for $\Omega>10^{1}$, JSf results in comparable distributions to that of the exact method.
Unlike in the previous example, here we observe that all JSF examples require more CPU time to complete the equivalent computation, despite this being a decaying death process. However, we still observe that all JSF methods require less CPU time when compared to the C++ implementation of the Tau-hybrid method, this time consuming at least 1.5 times less computation time, and still at least one order of magnitude quicker when compared to the Python implementation. We also note that JSF requires at most half the computational time of that of the exact Doob-Gillespie method. We now consider the time for the species to reach a population of size $X=50$. In this example, we see that JSF with $\Omega=10^{1}$ does not capture the distribution of the exact method, but only reports a single value. This is due to trajectories from this example all being deterministic at this value. If we consider JSF with $\Omega=10^{2}$, we see that while the median is comparable to the exact method, the distribution is not as broad. This is because the method only becomes stochastic at $X=100$, resulting in variance in the trajectory prior to this value being lost. As we increase the switching threshold to $\Omega=10^{3}$, we observe that distribution in the Time to reach $X=50$ become comparable to that of thee exact method. This provides a clear guide for how the switching threshold may be chosen, given some prior knowledge of the system. Lastly, we observe that extinction time for the JSF with $\Omega=10^{2}$ and $\Omega=10^{3}$ resulting in a comparable distribution to that of the Doob-Gillespie method. Again, we also observe that JSF with $\Omega=10^{1}$ results in a truncated distribution to that of the exact method.
These results also suggest that JSF is capable of recovering the key quantities of the exact stochastic method, while being computationally cheaper than current hybrid methods, for the death dominated birth-death process. Through this example, we also show how JSF may exhibit sensitivity in the value of the switching threshold chosen, and how to circumvent these sensitivities.
SI3 Simulation study: Multiple time-scales in autocatalysis
To emphasise the utility and accuracy of our Jump-Switch-Flow method, we will consider a simple autocatalysis example, with multiple time-scales. Consider two species, $X$, and $Y$, which react together via the following reactions: $X+Y\rightarrow 2X+Y$, with rate $\alpha$, $X\rightarrow 2X$, with rate $\beta$, $X\rightarrow\varnothing$, with rate $\gamma$, and $Y\rightarrow\varnothing$ with rate $\delta$. To demonstrate the ability of JSF to simulate the trajectories with multiple time-scales, but importantly, simulate them accurately, we choose the following rate parameters:
  1. 1.
    a moderate growth rate of $X$: $\beta=10$,
  2. 2.
    a large death rate of $Y$, to ensure we cannot assume a constant $Y$: $\delta=100$,
  3. 3.
    a large death rate of $X$, meaning $X$ quickly vanishes: $\gamma=50$,
  4. 4.
    and lastly a small autocatalysis rate, resulting in $X$ varying drastically in scale: $\alpha=1$.
Despite being a relatively simple model, with an analytic deterministic description, simulating stochastic trajectories requires careful consideration of the partitioning between the stochastic reactions to ensure an accurate representation, as there are reactions that are occurring a multiple different timescales. For example, autocatalysis occurs initially relatively frequently, due to the large number of species $Y$, despite the small rate $\alpha$. However, following the rapid decline of $Y$, $X$ will experience moderate decay. Therefore, we expect to observe two different timescales in the species $X$: a rapid increase, followed by gradual decline.
We assign the initial species sizes $X(0)=10$, and $Y(0)=10^{2}$, and sample 1000 trajectories using the exact Doob-Gillesie method, GillesPy2’s C++ hybrid method (Tau-hybrid) and JSF with a switching threshold of $\Omega=10^{2}$. Figure SI3 shows the 1000 trajectories for Tau-hybrid, JSF (with $\Omega=10^{2}$), and Doob-Gillespie. We also present the analytic ODE solutions (dashed black) and the median value at a point (solid black).
Figure SI3: 1000 trajectories for the multiscale autocatalyst example for the Tau-hybrid method (left), the JSF method (middle) and the exact Doob-Gillespie method (right). Top shows species $X$, and bottom shows species $Y$. We also show the median species value (solid black line) and the analytic ODE solution (dashed black line). We see that the median trajectory of Tau-hybrid is less than the analytic solution, while both Doob-Gillespie and JSF match the analytic solution.
Here, we can see that all $Y$ species solutions match the analytic solution, since species $Y$ undergoes constant decay only. However, we clearly see that, while the median trajectories of JSF and Doob-Gillespie match that of the analytic description, the Tau-hybrid method median is less than the analytic solution. We can further quantify these discrepancies by looking at the distributions of the time for species $X$ to reach a particular value, the distributions of the maximum $X$ value, and also the relative point error between the median of the particular approach, and the analytic solution. We present these quantities in Figure SI4.
Figure SI4: Summary statistics for the multiscale autocatalyst example for the Tau-hybrid (red), JSF with $\Omega=10^{3}$ (purple) and Doob-Gillespie (yellow). Left: time to reach $X=50$ following initial peak. We see that the distributions of Doob-Gillespie and JSF are comparable, while tau-hybrid is skewed downward with a different median value. Middle: Maximum $X$ value. Again, we observe JSF and Doob-Gillespie resulting in matching distributions, while Tau-hybrid is skewed downwards. Right: relative error to analytic solution. JSF and Doob-Gillespie display similarly small error, an order of magnitude less than Tau-hybrid.
Figure SI4 shows the distributions for time to reach $X=50$ following the peak, where we see that JSF and Doob-Gillespie produce comparable results. However we also observe that the Tau-hybrid method under-performs, and results in earlier hitting times. We observe a similar trend in in the maximum $X$ value, with Tau-hybrid producing a skewed downwards distribution, and JSF and Doob-Gillespie producing comparable distributions. Lastly, the relative point error of JSF and Doob-Gillespie are comparable, with Tau-hybrid an order of magnitude greater.
From the results presented here, we see that our Jump-Switch-Flow method is capable of producing highly accurate hybrid trajectories, even with highly multiscale systems, as the one presented here. We also observe the summary statistics produced by JSF being comparable to those of the exact Doob-Gillespie method, and outperforming the highly efficient and sophisticated Tau-hybrid method.
SI4 Simulation study: SIRS with demography
@@
We first present how the SIRS model with demography can be expressed with the Jump-Switch-Flow method. For the Jump-Switch-Flow method to be a useful method, it must reproduce behaviour that is representative of the exact process (i.e. the CTMC). That is, given some summary statistic, distributions produced via sampling the JSF process should be comparable to those obtained via sampling the CTMC. Moreover, we require that it does so with acceptable computational efficiency. We show, through simulation experiments, how the Jump-Switch-Flow method compares to the Doob-Gillespie method [16] for accuracy, and how our approach can exhibit sensitivity with respect to model inputs. We present how our Jump-Switch-Flow method compares to both the Doob-Gillespie and Tau-Leaping methods.
@@
Figure 5(c) shows the SIRS model with demography as a Jump-Switch-Flow process. In this example figure, the $S$ and $R$ compartments are continuous (flowing), while the $I$ compartment is discrete (jumping). To decide which of the arrows/reactions are jumping, i.e. in $\mathcal{S}$, and which are flowing, i.e. in $\mathcal{S}^{\prime}$, we consider the reactants and products of the arrows: if an arrow has any discrete reactant or product, then that arrow is also jumping, otherwise it is flowing. Table SI1 shows the reactants and products of the SIRS model with demography.
Figure SI5: 5(a) The labelled compartmental SIRS model with demography. 5(b) The SIRS model with demography where the arrows describe the interactions between the reactants and products. Here, blue arrows do not result in any flow through the system, and can be ignored. 5(c) The model as a Jump-Switch-Flow system with continuous (flowing) S and R (black), and discrete (jumping) I (purple). Since I is jumping, the reactions involving I are modelled as discrete, jumping reactions.
@@
Table SI1: The stoichiometric matrices for the SIRS model with demography. The (*) indicates reactions that are jumping when $I$ is discrete and $S$ and $R$ are continuous. The rows with blue coloured entries do not alter the exchange through the system, and so need not be represented.
For the remainder of this section, we suppose parameter values for the SIRS model with demography that would be typical for a newly introduced, yearly seasonal communicable disease, as presented in Table SI2. With these parameter values, on average, we expect an infected individual to infect 2 other people per week (during the initial outbreak), and be infectious for 1 week, with immunity to the disease lasting for 1 year. For the population turnover, we assume individuals live for, on average, 85 years. Initially there are two infectious individuals (i.e. $I(0)=2$), no recovered individuals (i.e. $R(0)=0$), and the remainder of the population is susceptible (i.e. $S(0)=N(0)-I(0)$).
@@
Value 2/72/72 / 7 1/71/71 / 7 1/3651/3651 / 365 1/(85×365)1/(85$\times$ 365)1 / ( 85 × 365 ) 1/(85×365)1/(85$\times$ 365)1 / ( 85 × 365 )
Table SI2: Parameter values used for SIRS model with demography.
@@
    Extinction: all of the individuals recover before the disease manages to spread significantly throughout the population (see Figure 6(a) and 6(d));
@@
    Fade-out: the disease spreads throughout the population, however it fades-out in the trough proceeding the first wave (see Figure 6(b) and 6(e));
@@
    Endemic: the disease persists indefinitely within the population (see Figure 6(c) and 6(f)).
@@
Figure SI6: Possible simulated trajectories of SIRS model with demography for an initial population size $N(0)=10^{5}$, from Doob-Gillespie (yellow) and Jump-Switch-Flow (purple, $\Omega=10^{3}$) methods: 6(a) two examples of extinction trajectories of the infectious compartment, with associated $S-I$ phase plane in 6(d); 6(b) two examples of fade-out trajectories of the infectious compartment, with associated $S-I$ phase plane in 6(e); 6(c) two examples of endemic trajectories of the infectious compartment, with associated $S-I$ phase plane in 6(f).
SI4.1 Simulation experiments: Comparing Jump-Switch-Flow and Doob-Gillespie
We compare our Jump-Switch-Flow method to the gold standard approach for simulating exact solutions to CTMCs, the Doob-Gillespie method. To do so, we first specified an initial population size of $N(0)=10^{5}$, and a switching threshold of each compartment of $\Omega=10^{3}$. We then generated 5,000 simulations using the parameters in Table SI2.
To capture the dynamics of the model, we track the infectious compartment as a key compartment of interest. In doing so, we then obtain distributions for the peak number of infectious individuals and also the time for the infection to peak, for the fade-out scenario. We also track the cumulative number of infected individuals after a certain amount of time for the endemic scenario. These results are presented in Figure SI8.
@@
Figure SI7: We compare the infectious compartment from realisations for Jump-Switch-Flow 7(b) and Doob-Gillespie 7(g) solutions, for a population size of $N(0)=10^{5}$. We use a switching threshold of $\Omega=10^{3}$. We also show the distributions of the time to infectious peak (7(a) and 7(f)), and infectious peak values (7(c) and 7(h)). We compare the cumulative infections over the simulations for Jump-Switch-Flow 7(d) and Doob-Gillespie 7(i) solutions and the final distributions (7(e) and 7(j)). Point averages are also presented as a solid line.
Comparing Figures 7(a) and 7(f), we can immediately observe that the distributions for time to peak for Jump-Switch-Flow method appears to be representative of that for the Doob-Gillespie method. This is because the delay in the epidemic take off at low population is fully captured with this sufficiently large choice of switching threshold, $\Omega=10^{3}$, enabling the full stochastic effects to be accurately represented.
However, if we compare the distributions for the peak number of infectious individuals (7(c) and 7(h)), we can see that they differ significantly, for this chosen switching threshold. This difference can be understood if we consider how the process is behaving. Once the infectious compartment is sufficiently large for the Jump-Switch-Flow method, the method switches from a stochastic representation to a deterministic representation. Therefore, once the infectious compartment becomes deterministic, given that the susceptible population is also deterministic, its peak value is determined by the number of both infectious and susceptible individuals at this time. Since the number of infectious people at this time will always be the same, $I=\Omega=10^{3}$ people, any variation in peak value is due to the number of susceptible individuals, $S$, at the time of switching.
If the summary statistic we instead compare is the distribution of the cumulative number of infected individuals after four years of epidemic circulation, we observe that, even for a relatively low switching threshold, the Jump-Switch-Flow method performs well comparatively to the Doob-Gillespie method, see Figures 7(e) and 7(j). We also note that for these parameter choices and switching threshold, the infectious compartment switches between jumping and flowing states multiple times, yet the Jump-Switch-Flow method produces a cumulative distribution comparable to that of Doob-Gillespie. This indicates that the Jump-Switch-Flow method is robust for long term simulations.
The final summary statistics we are interested in are the probabilities of extinction, fade-out and endemic scenarios. Table SI3 shows the probability of each scenario computed from 5,000 simulated samples. Here, we see that both methods produce comparatively similar results, which indicates that the Jump-Switch-Flow method is capable of capturing the inherent stochasticity in the exact simulation of the system via the Doob-Gillespie method.
@@
Jump-Switch-Flow 0.2476±0.01200.2476$\pm$ 0.01200.2476 ± 0.0120 0.4774±0.01380.4774$\pm$ 0.01380.4774 ± 0.0138 0.2750±0.01240.2750$\pm$ 0.01240.2750 ± 0.0124
Doob-Gillespie 0.2520±0.01200.2520$\pm$ 0.01200.2520 ± 0.0120 0.4668±0.01380.4668$\pm$ 0.01380.4668 ± 0.0138 0.2812±0.01250.2812$\pm$ 0.01250.2812 ± 0.0125
Table SI3: Probabilities (with associated 95% confidence interval) of scenario outcomes for Jump-Switch-Flow and Doob-Gillespie, from 5,000 simulations.
We also compare the Jump-Switch-Flow and Doob-Gillespie methods where the switching threshold has been set to the population size, resulting in no compartments switching into the flowing state. These results are further discussed in SI4.2, where we show the two methods are indistinguishable.
SI4.2 Simulation experiments: Comparing Jump-Switch-Flow (with no switching) and Doob-Gillespie
We compare the Jump-Switch-Flow method to the Doob-Gillespie method. To implement the latter, we set the switching threshold to be the total population size, to ensure no compartments switch into a flowing regime, and therefore all compartments are completely stochastic. We can therefore investigate how our Jumping process compares to exact solutions of the Continuous Time Markov Chains.
@@
Figure SI8: We compare the infectious compartment from realisations for Jump-Switch-Flow 8(b) and Doob-Gillespie 8(g) solutions, for a population size of $N_{0}=10^{5}$. We use a switching threshold of $\Omega=10^{3}$. We also show the distributions of the time to infectious peak (8(a) and 8(f)), and infectious peak values (8(c) and 8(h)). We compare the cumulative infections over the simulations for Jump-Switch-Flow 8(d) and Doob-Gillespie 8(i) solutions and the final distributions (8(e) and 8(j)). Point averages are also presented as a solid line.
@@
Table SI4 shows that the probabilities of scenario outcomes are also comparable for both Jump-Switch-Flow (with no switching) and Doob-Gillespie.
@@
Jump-Switch-Flow 0.2512±0.01200.2512$\pm$ 0.01200.2512 ± 0.0120 0.4700±0.01390.4700$\pm$ $0$ 0.0139 0.2788±0.01230.2788$\pm$ 0.01230.2788 ± 0.0123
Doob-Gillespie 0.2520±0.01200.2520$\pm$ 0.01200.2520 ± 0.0120 0.4668±0.01390.4668$\pm$ 0.01390.4668 ± 0.0139 0.2812±0.01250.2812$\pm$ 0.01250.2812 ± 0.0125
Table SI4: Probabilities (with associated 95% confidence interval) of scenario outcomes for Jump-Switch-Flow (with no switching) and Doob-Gillespie, from 5,000 simulations.
SI4.3 Simulation experiments: Sensitivity of Jump-Switch-Flow to switching threshold
In this section, we demonstrate how the accuracy of the Jump-Switch-Flow method can exhibit sensitivity with respect to the switching threshold (Figure SI9). In Figure 9(a), we can see that the distribution of peak number of infective individuals for the fade-out scenario is highly sensitive to the switching threshold. For a small switching threshold ($\Omega=10^{1}$), we see that the distribution exhibits low variance, is representative of the deterministic expectation, and is not capturing the distribution obtained with the Doob-Gillespie method. As the switching threshold is increased, we see that the distribution of the peak number of infective individuals increases in variance. However, the Jump-Switch-Flow method is not representative of the Doob-Gillespie method until a large switching threshold ($\Omega=N(0)=10^{5}$), where all compartments are in the jumping state.
Figure SI9: Summary statistics of interest to observe the sensitivity of the Jump-Switch-Flow method with respect to the switching threshold. 9(a) Sensitivity to the distribution of peak number of infective individuals for the fade-out scenario. 9(c) Sensitivity to the distribution of peak infection time for the fade-out scenario. 9(b) Sensitivity to the distribution of the cumulative number of infected individuals after four years for the endemic scenario. 9(d) Sensitivity to the probability of each scenario occurring. In each of the plots 9(a)-9(c) grey to purple shaded distributions represent results obtained with the Jump-Switch-Flow method whilst the yellow distributions are the gold standard Doob-Gillespie algorithm (for which the placeholder ‘G’ is used to signify) which exhibit exact stochasticity but at the cost of high computational requirements.
@@
SI4.4 Simulation experiments: Computational efficiency of Jump-Switch-Flow
@@
We now present the computational efficiency of the Jump-Switch-Flow method, and compare it to the Doob-Gillespie method and Tau-Leaping method (which is regarded as a computationally efficient, approximate, method), and the Tau-hybrid method, provided by GillesPy2 [33]. We implement two variants of Tau-leaping with a fixed step size of $\Delta\tau=0.01$ days, and $\Delta\tau=0.1$ days respectively, two variants of the C++ implementation of the Tau-hybrid method, with time steps of size $\Delta t=0.01$ days and $\Delta t=0.1$ days respectively, one implementation of the Python Tau-hybrid method, with time steps of size $\Delta t=0.1$ days, and the Jump-Switch-Flow method with a time step size of $\Delta t=0.01$ days.